OpenStudy (oswaldmurphy):

The sides of a square are five to the power of two fifths inches long. What is the area of the square? five to the power of four fifths square inches five to the power of four twenty fifths square inches twenty five to the power of four fifths square inches twenty five to the power of four twenty fifths square inches I got D is that correct?

TheSmartOne (thesmartone):

nope :P Hint: \(\Large\sf a^b \times a^c = a^{b+c}\)

OpenStudy (oswaldmurphy):

I didnt realize that was a power..

OpenStudy (oswaldmurphy):

A?

TheSmartOne (thesmartone):

\(\Large\sf a^{\frac{b}{d}} \times a^{\frac{c}{d}} = a^{\frac{b+c}{d}}\)

TheSmartOne (thesmartone):

correct

OpenStudy (oswaldmurphy):

Ok. Thanks bb

OpenStudy (jtug6):

Just to add as an additional rule as reference (also important not to confuse with the multiplication property) \[(A ^{7})^{3} = A ^{21}\]

OpenStudy (oswaldmurphy):

OK, thanks :D

OpenStudy (jtug6):

Raising a power to another power = multiplication

OpenStudy (jtug6):

np